# Non-linear anomalous Edelstein response at altermagnetic interfaces

### v2

@@ Abstract
In altermagnets, time-reversal symmetry breaking spin-polarizes electronic states, while total magnetization remains zero. In addition, at altermagnetic surfaces Rashba-spin orbit coupling is activated due to broken inversion symmetry, introducing a competing spin-momentum locking interaction. Here we show that their interplay leads to the formation of complex, chiral spin textures that offer novel, non-linear spin-to-charge conversion properties. Whereas altermagnetic order suppresses the canonical linear in-plane Rashba-Edelstein response, we establish the presence of an anomalous transversal Edelstein effect for planar applied electric and magnetic field, or alternatively, an in-plane magnetization. Moreover the non-linear Edelstein response resulting purely from electric fields also triggers the anomalous out-of-plane magnetization. We determine the anomalous response with a model based on the ab-initio electronic structure of RuO2 bilayers, ultimately opening experimental avenues to explore spin-charge conversion phenomena at altermagnetic interfaces.
Altermagnets (AMs) are a class of magnetic materials with a total magnetization that vanishes by symmetry while at the same time Kramers’ degeneracy is lifted [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 3, 11]. Unlike antiferromagnets, for which the vanishing of the magnetization is enforced by the combined action of time-reversal and translation/inversion symmetry, in AMs this involves an additional rotation or mirror operation. This symmetry operation causes the band structure to be spin-polarized, also in the absence of spin-orbit coupling (SOC) [1]. The presence of spin-polarized bands and Fermi surfaces makes AMs potentially appealing as spin-polarized current sources for antiferromagnetic spintronics [5, 12, 13, 14, 15, 16]. On the basis of magnetic symmetries [1], an extensive set of candidate materials exhibiting collinear altermagnetism have been identified from electronic structure calculations, both 3D [8] and 2D [17]. While experimentally the spin-splitting of electronic bands has been observed by photoemission spectroscopy [18, 19, 20, 21, 22], exploration of AM (magneto–) transport properties for antiferromagnetic spintronic purposes, which capitalizes on their strong exchange interactions [23], has only just begun and has so far focused on harnessing their anomalous and spin Hall effects [24, 9, 13, 25, 26, 15, 19, 27]. For instance it has recently been shown that altermagnetic spin-splitting is at the origin of a spin Hall effect in RuO2 [12].
@@
Figure 1: Schematic representation of the transversal, non-linear altermagnetic EE. (a) Out-of-plane spin arrangement in the absence of driving current with zero net magnetization and (b) homogeneous tilting of the out-of-plane spins due to an in-plane driving current that generates a transverse magnetic field.
In this context, we identify natural possibilities for spin-to-charge conversion processes at AM interfaces that fundamentally go beyond the canonical linear Rashba-Edelstein effect [28, 29, 30], where a surface electric current produces an in-plane magnetization that, as we will show, is actually suppressed by AM. Instead, the competition between Rashba SOC (RSOC) and AM triggers anomalous non-linear Edelstein effects (EEs): we will show that a magnetization transverse to the interface is induced by an in-plane applied electric and magnetic field (or an in-plane magnetization) – or alternatively as non-linear magnetic response to purely planar electric fields. This spin-to-charge conversion in electric and magnetic field can be depicted by an in-plane current forcing the magnetic moments out of the plane, as schematically indicated in Fig. 1. Also the purely electric non-linear case will be understood in such basic terms.
@@
It should be noted that, while AM may be strictly defined in the limit of vanishing SOC, it still defines an essential energy scale, determining for instance the altermagnetic anomalous Hall response [9] and electronic properties related to the topology of altermagnetic electronic structures [24]. At interfaces and in 2D, the RSOC is induced by inversion-symmetry breaking and naturally produces a spin-momentum locking without net magnetization. Indeed, the coexistence of AM spin splitting and the RSOC has been pointed out [31, 32], but so far not investigated as a mechanism to achieve an anomalous Rashba-Edelstein response and non-linear spin-to-charge conversion.
@@
Figure 2: Competing altermagnetic and Rashba spin-orbit interactions. (a) Electronic dispersion resulting from the continuum Hamiltonian (1) $\alpha=52$ meV Å. (b) Fermi surface and spin structure for fixed chemical potential (dashed line in panel (a)). (c) linear Edelstein susceptibility $\chi_{xy}/\chi_{0}$ ($\chi_{0}=2.44\times 10^{2}$ $\mu_{B}$ V-1) for different $\alpha$. The dashed lines correspond to the pure RSOC case ($\gamma=0$).
@@
The standard linear EE originating from RSOC locks the spin of electrons in-plane and orthogonal to their momentum, respecting spinfull time-reversal symmetry. Then, an applied in-plane electric field results in an imbalance in the spin density at the Fermi surface and a net in-plane magnetization orthogonal to the induced current. Interestingly, in the presence of only RSOC, the second-order non-linear EE [33, 34, 35] vanishes by symmetry. However, it may be triggered by, for instance, introducing orbital degrees of freedom [33] or a time-dependent setting [34]. In contrast to RSOC, AM breaks time-reversal symmetry, so that in collinear AMs opposite momenta carry the same spin orientation. This renders the momentum-space magnetic textures generated by RSOC and AM fundamentally different: the two interactions compete.
@@
Altermagnetic Rashba model — To explore the physical consequences of this competition, we start by considering the 2D effective continuum Hamiltonian
@@
where $\hat{\sigma}_{i}$ is the Pauli matrix in the direction $\hat{i}$, $\alpha$ is the RSOC interaction, $\gamma$ is the altermagnetic $d$-wave coupling with symmetry $[C_{2}||C_{4z}]$ [1], $\mu$ is the chemical potential, and $\hbar$ is set to unity. The resulting bands and their spin-texture for given chemical potential are illustrated in Fig. 2(a,b). Note that the chirality of the resulting non-coplanar spin texture is determined by $\mathrm{sgn}(\gamma\alpha)$.
To keep contact with real materials, we consider a bilayer of RuO2 and estimate its altermagnetic coupling $\gamma$. We note that the magnetic nature of RuO2 is in itself an interesting debated question [36, 37, 38, 39, 40]. Here, we consider this system just as a $d$-wave altermagnetic model in which it is in principle possible to engineer via gating both the band filling and effective RSOC. Based on the bulk crystal structure, we construct a bilayer oriented along the $c$-axis of RuO2 bulk, containing two Ru ions related by fourfold rotoinversion symmetry. The altermagnetic bilayer can be represented with the space group Cmm2 (No. 35), with a unit cell having $a=6.35$. For this structure, we perform density-functional calculations in the generalized gradient approximation based on the FPLO code [41]. We use a $k$-points mesh having $16\times 16\times 1$ subdivisions along with a tetrahedron method for Brillouin zone integrations. Figure 3 shows the band structure in the absence of RSOC. We apply our effective two-band model to describe the two bands nearest to the Fermi energy at $\Gamma$. From a fit of Eq. (1) to these bands, we obtain $m=0.152$ eV-1 Å-2 and $\gamma=1.881$ eV Å2 (see Appendix B).
Figure 3: Bilayer of RuO2 and its electronic structure density-functional calculations. (a) Crystal structure. (b) Band structure along high symmetry points indicated in (c).
Linear Edelstein responses — The linear EE is defined through $\expectationvalue{\sigma_{j}}=\chi_{ij}E_{i}$, where $E_{i}$ is the electric field along the direction $\hat{i}$ and $\chi_{ij}$ is the Edelstein susceptibility. The spin accumulation $\expectationvalue{\sigma_{j}}$ can be obtained in the Boltzmann framework as
$\expectationvalue{\sigma_{j}}(\mu)=-\mu_{b}\sum_{\mathbf{k},\nu}|e|\left(\vec{\Lambda}_{\mathbf{k}}^{\nu}\cdot\mathbf{E}\right)\delta\left[\varepsilon_{\mathbf{k}}^{\nu}-\mu\right](\nu s_{j}(\mathbf{k})),$ (2)
@@
Supposing, for simplicity, that $\tau^{\nu}_{\mathbf{k}}=\tau$, and taking $\mathbf{E}=E_{0}\hat{x}$, Eq. (2) can be written in terms of the susceptibility per unit cell, $\chi_{xi}$, as
$\chi_{xj}=-\chi_{0}\sum_{\nu=\pm}\int d^{2}\mathbf{k}\;\nu s_{j}(\mathbf{k})\delta(\varepsilon_{\mathbf{k}}^{\nu}-\mu)v^{\nu}_{x}(\mathbf{k}),$ (4)
@@
The conventional EE is governed by $\chi_{xy}$ and certainly RSOC alone is sufficient to induce it, as shown for our continuum Hamiltonian in Fig 2(c). Without AM and in-plane magnetic field ($\gamma=0$), the linear Edelstein susceptibility $\chi_{xy}$ scales with the RSOC $\alpha$, and quickly saturates for large chemical potential $\mu$, see dashed lines in Fig 2(c). Here, we vary $\alpha$ in a range around $0.05$ eV Å which is a realistic range of values for oxides [42, 43, 44, 45], and may be experimentally achieved by, e.g., a tunable top-gate or interfacial electric field [46]. Clearly, a finite altermagnetic coupling, $\gamma$, significantly suppresses the EE, particularly at higher $\mu$, and this for any value of $\gamma$ leads to a pronounced peak in $\chi_{xy}$ in the vicinity of the band bottom. Thus, as we expected on the basis of symmetry considerations, AM and the conventional EE strongly compete.
@@
Definition of non-linear Edelstein responses — Clearly, the RSOC, with its in-plane spin-momentum locking, cannot alone produce an out-of-plane, transverse Edelstein response, i.e., $\chi_{xz}$ vanishes identically. Even if the $\gamma$ term tends to align spins out-of-plane, RSOC in the presence of altermagnetism is not sufficient to induce a transversal response. Since $\varepsilon^{\pm}_{\mathbf{k}}=\varepsilon^{\pm}_{-\mathbf{k}}$, while $v_{x}^{\pm}(\mathbf{k})=-v_{x}^{\pm}(-\mathbf{k})$, the component $\chi_{xz}$ is seen to vanish being $s_{z}(-\mathbf{k})=s_{z}(\mathbf{k})$. Such an anomalous transversal response actually requires altermagnetism and a non-linear EE, as we will detail in the following. Indeed, the presence of an additional (effective) in-plane magnetic field breaks the $\mathbf{k}\to-\mathbf{k}$ symmetry, thus allowing out-of-plane Edelstein responses. One may distinguish two types of non-linear response, one involving an in-plane electric and magnetic field and one involving only in-plane electric fields, but up to second order (see Appendix A).
@@
The case of combined electric and magnetic fields is most easily considered by introducing a Zeeman term in Eq. (1), orthogonal to the applied electric field $E_{x}$, of the form $H_{h}=h_{y}\sigma_{y}$ and then using Eq. (4). The $H_{h}$ term corresponds to an effective homogeneous spin splitting, which may arise from a Zeeman field or from weak ferromagnetism induced by Dzyaloshinskii–Moriya (DM) interactions, generally allowed by AM symmetries [47, 17, 48, 49]. This response is nonlinear as it depends both linearly on electric field and magnetic coupling and vanishes when either one is zero (see Appendix C). Incorporating $\mathbf{h}$ directly into the Hamiltonian offers the advantage that it is easy to go beyond its linear response.
@@
To assess the non-linear EE involving only electric fields, it is more convenient to introduce the second order Edelstein susceptibility of the form $\expectationvalue{\sigma_{z}}=\tilde{\chi}_{xxz}E_{x}^{2}$, with
$\tilde{\chi}_{xxz}=-\tilde{\chi}_{0}\sum_{\nu=\pm}\int d^{2}\mathbf{k}\;\nu\frac{\partial s_{z}(\mathbf{k})}{\partial k_{x}}\delta(\varepsilon^{\nu}_{\mathbf{k}}-\mu)v_{x}^{\nu}(\mathbf{k}),$ (5)
where $\tilde{\chi}_{0}=\frac{\mu_{B}e^{2}\tau^{2}}{4\pi^{2}}S_{\text{cell}}$ (see Appendix D for the full derivation). In this case, the symmetry argument discussed above shows that the integrand function is even rather than odd, since $\partial_{k_{x}}s_{z}(\mathbf{k})=-\partial_{k_{x}}s_{z}(-\mathbf{k})$. Therefore, there are no symmetries that force the response to vanish.
@@
Figure 4: Anomalous non-linear EE due to a combined planar magnetic and electric field. (a) Electronic dispersion from Hamiltonian (1) with $\alpha=52$ meV Å and $h_{y}=40$ meV. (b) Fermi surface and spin structure for fixed chemical potential, see the dashed line in (a). (c) Edelstein response $\chi_{xi}/\chi_{0}$ ($\chi_{0}=2.44\times 10^{2}$ $\mu_{B}$ V-1) for the directions $i={x,y,z}$. The green dashed line corresponds to the pure RSOC case ($\gamma=h_{y}=0$).

### v3

@@ Abstract
In altermagnets, time-reversal symmetry breaking spin-polarizes electronic states, while total magnetization remains zero. In addition, at altermagnetic surfaces Rashba-spin orbit coupling is activated due to broken inversion symmetry, introducing a competing spin-momentum locking interaction. Here we show that their interplay leads to the formation of complex, chiral spin textures that offer novel, non-linear spin-to-charge conversion properties. Whereas altermagnetic order suppresses the canonical linear in-plane Rashba-Edelstein response, we establish the presence of an anomalous transversal Edelstein effect for planar applied electric and magnetic field, or alternatively, an in-plane magnetization. Additionally, we predict a purely electric-field-driven non-linear out-of-plane magnetization. We compute the anomalous response within a general altermagnet $d$-wave model, with parameters extracted from the ab-initio electronic structure of an altermagnetic bilayer. Our results suggest altermagnetic surfaces as a promising platform for unconventional spintronic functionalities.
I Introduction
@@
Figure 1: Schematic representation of the transversal, non-linear altermagnetic $E$. (a) Out-of-plane spin arrangement in the absence of driving current with zero net magnetization. (b) An in-plane driving current tilts the out-of-plane spins resulting in an in-plane transverse magnetic field. (c) A stronger current is able to further tilt the spins resulting in a non-linear magnetization, which is also out-of-plane.
Altermagnets (AMs) are a class of magnetic materials with a total magnetization that vanishes by symmetry while at the same time Kramers’ degeneracy is lifted [45, 37, 1, 43, 39, 57, 40, 19, 42, 44, 1, 36]. Unlike antiferromagnets, for which the vanishing of the magnetization is enforced by the combined action of time-reversal and translation/inversion symmetry, in AMs this involves an additional rotation or mirror operation. This symmetry operation causes the band structure to be spin-polarized, also in the absence of spin-orbit coupling (SOC) [45]. The presence of spin-polarized bands and Fermi surfaces makes AMs potentially appealing as spin-polarized current sources for antiferromagnetic spintronics [39, 6, 18, 58, 5, 7, 16]. On the basis of magnetic symmetries [45], an extensive set of candidate materials exhibiting collinear altermagnetism have been identified from electronic structure calculations, both 3D [19] and 2D [38]. While experimentally the spin-splitting of electronic bands has been observed by photoemission spectroscopy [59, 41, 35, 24, 61], exploration of AM (magneto–) transport properties for antiferromagnetic spintronic purposes, which capitalizes on their strong exchange interactions [12], has only just begun and has so far focused on harnessing their anomalous and spin Hall effects [31, 42, 18, 15, 17, 5, 41, 21]. For instance it has recently been shown that altermagnetic spin-splitting is at the origin of a spin Hall effect in RuO2 [6]. In this context, we identify natural possibilities for spin-to-charge conversion processes at AM interfaces that fundamentally go beyond the canonical linear Rashba-Edelstein effect [13, 50, 51], namely the generation of an in-plane magnetization in response to a planar electric field caused by antisymmetric spin-texture at the Fermi surface. Indeed, we show that this response is actually suppressed by AM. Instead, the competition between Rashba SOC (RSOC) and AM triggers anomalous non-linear Edelstein effects (EEs): we will show that a magnetization transverse to the interface is induced by an in-plane applied electric and magnetic field (or an in-plane magnetization) – or alternatively as non-linear magnetic response to purely planar electric fields. This spin-to-charge conversion in electric and magnetic field can be depicted by an in-plane current forcing the magnetic moments out of the plane, as schematically indicated in Fig. 1. Also the purely electric non-linear case will be understood in such basic terms.
@@
It should be noted that, while AM may be strictly defined in the limit of vanishing SOC, it still defines an essential energy scale, determining for instance the altermagnetic anomalous Hall response [42] and electronic properties related to the topology of altermagnetic electronic structures [31]. At interfaces and in 2D, the RSOC is induced by inversion-symmetry breaking and naturally produces a spin-momentum locking without net magnetization. Indeed, the coexistence of AM spin splitting and the RSOC has been pointed out [47, 2, 10], but so far not investigated as a mechanism to achieve an anomalous Rashba-Edelstein response and non-linear spin-to-charge conversion.
@@
Figure 2: Competing altermagnetic and Rashba spin-orbit interactions. (a) Electronic dispersion resulting from the continuum Hamiltonian (1) $\alpha=52$ meV Å. (b) Fermi surface and spin structure for fixed chemical potential (dashed line in panel (a)). (c) linear Edelstein susceptibility $\chi_{xy}/\chi_{0}$ ($\chi_{0}=2.44\times 10^{2}$ $\mu_{B}$ Å V-1) for different $\alpha$. The dashed lines correspond to the pure RSOC case ($\gamma=0$).
@@
The standard linear EE originating from RSOC locks the spin of electrons in-plane and orthogonal to their momentum, respecting spinful time-reversal symmetry. Then, an applied in-plane electric field results in an imbalance in the spin density at the Fermi surface and a net in-plane magnetization orthogonal to the induced current. Interestingly, in the presence of only RSOC, the second-order non-linear EE [4, 52, 56] vanishes by symmetry. However, it may be triggered by, for instance, introducing orbital degrees of freedom [4] or a time-dependent setting [52]. In contrast to RSOC, AM breaks time-reversal symmetry, so that in collinear AMs opposite momenta carry the same spin orientation. This renders the momentum-space magnetic textures generated by RSOC and AM fundamentally different: the two interactions compete.
II Altermagnetic Rashba model
@@
To explore the physical consequences of this competition, we start by considering the 2D effective continuum Hamiltonian
@@
where $\hat{\sigma}_{i}$ is the Pauli matrix in the direction $\hat{i}$, $\alpha$ is the RSOC interaction, $\gamma$ is the altermagnetic $d$-wave coupling with symmetry $C_{2}||C_{4z}] [$45], $\mu$ is the chemical potential, and $\hbar$ is set to unity. The resulting bands and their spin-texture for given chemical potential are illustrated in Fig. 2(a,b). Note that the chirality of the resulting non-coplanar spin texture is determined by $\mathrm{sgn}(\gamma\alpha)$. To maintain contact with real materials, we choose the numerical values of the altermagnetic coupling $\gamma$ and the effective mass $m$ by explicitly comparing our $\bf k\cdot p$ model with the local band structure close to the $\Gamma$ point of a RuO2 bilayer. Details of ab-initio calculation and band-structure fitting are given in the Appendix B. We use $m=0.152$ eV-1 Å-2 and $\gamma=1.881$ eV Å2. While the magnetic nature of RuO2 is a debated question [33, 34, 22, 27, 54], thin films of RuO2 [8, 46, 20, 28, 14] have been characterized as $d$-wave altermagnets. Here, we consider this system just as a representative $d$-wave altermagnetic model , in which both the band filling and the effective RSOC can, in principle, be tuned via gating.
Linear Edelstein responses
The linear EE is defined through $M_{j}=\chi_{ij}E_{i}$, where $E_{i}$ is the electric field along the direction $\hat{i}$ and $\chi_{ij}$ is the Edelstein susceptibility. The spin accumulation $M_{j}$ can be obtained in the Boltzmann framework as
$M_{j}(\mu)=-\mu_{b}\sum_{\mathbf{k},\nu}|e|\left(\vec{\Lambda}_{\mathbf{k}}^{\nu}\cdot\mathbf{E}\right)\delta\left[\varepsilon_{\mathbf{k}}^{\nu}-\mu\right](\nu s_{j}(\mathbf{k})),$ (2)
@@
Supposing, for simplicity, that $\tau^{\nu}_{\mathbf{k}}=\tau$, and taking $\mathbf{E}=E_{0}\hat{i}$, Eq. (2) can be written in terms of the susceptibility per unit cell, $\chi_{ij}$, as
$\chi_{ij}=-a\chi_{0}\sum_{\nu=\pm}\int d^{2}\mathbf{k}\;\nu s_{j}(\mathbf{k})\delta(\varepsilon_{\mathbf{k}}^{\nu}-\mu)v^{\nu}_{i}(\mathbf{k}),$ (4)
@@
The conventional EE is governed by $\chi_{xy}$ and RSOC alone is sufficient to induce it, as shown for our continuum Hamiltonian in Fig 2(c). The symmetry group of the Hamiltonian $C_{2}||C_{4z}]$ with broken inversion symmetry enforces $\chi_{xx}=\chi_{yy}=0$, $\chi_{xy}=-\chi_{yx}$, and $\chi_{iz}=0$ for $i=x,y$. Without AM and in-plane magnetic field ($\gamma=0$), the linear Edelstein susceptibility $\chi_{xy}$ scales with the RSOC $\alpha$, and quickly saturates for large chemical potential $\mu$, see dashed lines in Fig 2(c). Here, we vary $\alpha$ in a range around $0.05$ eV Å which is a realistic range of values for oxides [30, 48, 49, 60], and may be experimentally achieved by, e.g., a tunable top-gate or interfacial electric field [9]. Clearly, a finite altermagnetic coupling, $\gamma$, significantly suppresses the EE, particularly at higher $\mu$, and this for any value of $\gamma$ leads to a pronounced peak in $\chi_{xy}$ in the vicinity of the band bottom. Thus, as we expected on the basis of symmetry considerations, AM and the conventional EE strongly compete.
Definition of non-linear Edelstein responses
@@
Clearly, the RSOC, with its in-plane spin-momentum locking, cannot alone produce an out-of-plane, transverse Edelstein response, i.e., $\chi_{xz}$ vanishes identically. Even if the $\gamma$ term tends to align spins out-of-plane, RSOC in the presence of altermagnetism is not sufficient to induce a transversal response. Such an anomalous transversal response actually requires altermagnetism and higher orders in the fields, as we will detail in the following. One may distinguish two types of non-linear (second order) responses, one involving an in-plane electric and magnetic field and one involving only in-plane electric fields to second order. In the following, we focus solely on the extrinsic contribution arising from impurity scattering. For such a model, the intrinsic interband Edelstein contribution, originating from the off-diagonal terms [32], vanishes by symmetry (see details in Appendix F). Other intrinsic effects, such as those related to quantum geometry [55] or deviations from the rigid-band approximation [23, 26], could in principle be considered, but they are expected to be more relevant for complex multiband structures. Here, we highlight the standard contribution, which should also dominate for thin films on a substrate, where the constant-scattering-time approximation is valid [25].
@@
The case of combined electric and magnetic fields is most easily considered by introducing a Zeeman term in Eq. (1), orthogonal to the applied electric field $E_{x}$, of the form $H_{h}=h_{y}\sigma_{y}$ and then using Eq. (4). The $H_{h}$ term corresponds to an effective homogeneous spin splitting, which may arise from a Zeeman field or from weak ferromagnetism induced by Dzyaloshinskii–Moriya (DM) interactions, generally allowed by AM symmetries [53, 38, 11, 3]. The presence of the in-plane magnetic field spontaneously breaks the $\mathcal{T}C_{4z}$ symmetry, allowing out-of-plane Edelstein response. This response is nonlinear as it depends both linearly on electric field and magnetic coupling and vanishes when either one is zero (see Appendix A). Incorporating $\mathbf{h}$ directly into the Hamiltonian offers the advantage that it is easy to go beyond its linear response.
@@
To assess the non-linear EE involving only electric fields, it is more convenient to introduce the second order Edelstein susceptibility of the form $M_{j}=\tilde{\chi}_{iij}E_{i}^{2}$, with
$\tilde{\chi}_{iij}=-\tilde{\chi}_{0}\sum_{\nu=\pm}\int d^{2}\mathbf{k}\;\nu\frac{\partial s_{j}(\mathbf{k})}{\partial k_{i}}\delta(\varepsilon^{\nu}_{\mathbf{k}}-\mu)v_{i}^{\nu}(\mathbf{k}),$ (5)
where $\tilde{\chi}_{0}=\frac{\mu_{B}e^{2}\tau^{2}}{4\pi^{2}}S_{\text{cell}}$ (see Appendix D for the full derivation). The $\mathcal{T}C_{4z}$ symmetry does not rule out an out-of-plane response and enforces $\tilde{\chi}_{xxz}=-\tilde{\chi}_{yyz}$, (while it easily excludes an in-plane spin response $\tilde{\chi}_{iij}=0$ for $i,j=x,y$).
@@
Figure 3: Anomalous non-linear EE due to a combined planar magnetic and electric field. (a) Electronic dispersion from Hamiltonian (1) with $\alpha=52$ meV Å and $h_{y}=40$ meV. (b) Fermi surface and spin structure for fixed chemical potential, see the dashed line in (a). (c) Edelstein response $\chi_{xi}/\chi_{0}$ ($\chi_{0}=2.44\times 10^{2}$ $\mu_{B}$ Å V-1) for the directions i=x,y,zi={x,y,z}. The green dashed line corresponds to the pure RSOC case ($\gamma=h_{y}=0$).
III Calculation of non-linear Edelstein responses